OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of ten dollars for dinners that cost nine dollars, eight dollars, eleven dollars and twelve dollars each. Use a percent proportion to solve for the percent tip they left the waiter

OpenStudy (texaschic101):

9 + 8 + 11 + 12 = 40 -- total of all dinners 10 - tip left Basically you need to know : 10 is what percent of 40 percent proportion is/of = %/100 10/40 = x/100 cross multiply 40x = (10)(100) 40x = 1000 x = 1000/40 x = 25%

OpenStudy (imstuck):

Well you first have to add up the dinners to see what they total. That's 9 + 8 + 11 + 12 = 40. If they tipped him $10 and you need an equation to solve for the percent, you need to ask yourself, "what percent of 40 is 10?" The equation in numbers is\[10=x \times40\]That reads, "$10 is ("is" means equals) x (or "what percent") of ("of" means multiplication) $40" or 10 = 40x. Now solve for x. This is a small number, yes, but don't forget that we are converting to a percent, which is based on 100. So you need to multiply your answer by 100 to find the percentage.